OpenStudy (anonymous):

Which of the following equations represents a line parallel to the graph of 3x-y=7? y=7x-3 y=-3x+7 y=-7x+3 y=3x

OpenStudy (jdoe0001):

if you solve " 3x-y=7 " for "y" what would it give you ?

OpenStudy (ccswims):

Well "parallel" means they will have the same slope so 3x

OpenStudy (anonymous):

first you need to put in in slop form y=mx+b do you know how to do that?

OpenStudy (ccswims):

Oh wow, I didn't even see that, so yea it's not 3x.. It's the opposite

OpenStudy (anonymous):

it would be y=-3x+7

OpenStudy (anonymous):

the slope is -3 and the y intercept is 7

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

I think it would be y = -7 + 3x? @alenajonesy

OpenStudy (ccswims):

y=mx + b the y-intercept would be in the back

OpenStudy (anonymous):

no because it has to be in y=mx+b

OpenStudy (anonymous):

y= 3x+7 @alenajonesy

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait no -3x

OpenStudy (anonymous):

because your subtracting 3x from the y so it would be -3x

OpenStudy (anonymous):

I figured that part out when i went to the answers. lol.

OpenStudy (anonymous):

haha ok so anything else?

OpenStudy (anonymous):

There's a bunch of questions, if you wouldn't mind helping me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Okay. Plans for a bridge are drawn on a coordinate grid. One girder of the bridge lies on the line y=-9x-9. A perpendicular brace passes through the point (-1,7) Write an equation of the line that contains the brace.

OpenStudy (anonymous):

y-1=-1/9(x+7) x-7=-9(y+1) y-7=1/9(x+1) y-7=-9(x+1)

OpenStudy (anonymous):

the equation for this would be y-y1=m(x-x1)

OpenStudy (anonymous):

since its already in y=mx+b you don't have to change that you just need to plug in the coordinates and the slope

OpenStudy (anonymous):

So the answer is y - 7 = (1/9)(x + 1) ?

OpenStudy (anonymous):



OpenStudy (anonymous):

do you understand that or need me to redo a step?

OpenStudy (anonymous):

I understand.

OpenStudy (anonymous):

any others?